# Tailoring Dynamical Codes for Biased Noise: The X3Z3 Floquet Code

### v1

@@ III X3Z3 Bias-tailored Floquet Code
We now describe the X3Z3 bias-tailored Floquet code, which is shown in Figs. 1(b) and 1(c). This code is related to the CSS Floquet code by Hadamard gates applied to the qubits in strips [shown in grey in Fig. 1(b)] along vertical non-trivial cycles of the lattice, and is a Floquetified version of the domain wall colour code [36]. We therefore still have two types of plaquettes and edges, one type originating from Pauli $X$ plaquettes and edges before the Hadamard gates are applied, and the other originating from Pauli $Z$ operators. We refer to these modified operators as A-type and B-type, respectively. They are defined in Fig. 1(b). The measurement sequence is analogous to the CSS Floquet code sequence: $r\text{A}\rightarrow g\text{B}\rightarrow b\text{A}\rightarrow r\text{B}\rightarrow g\text{A}\rightarrow b\text{B}$, where $c\text{A}$ represents the measurement of A-type check operators along $c$-coloured edges, and similarly for $c\text{B}$. Just as with the CSS Floquet code [11], this code can also be defined on a planar lattice with boundary.
The decoding of the X3Z3 Floquet code under $Z$-biased noise is simplified by the presence of a symmetry in the decoding graphs. This is a space-time analogue of the symmetries present in bias-tailored static codes such as the XZZX code [35] and the domain wall colour code [36]. In such codes, under the code-capacity noise model (with only single-qubit Pauli errors), syndromes resulting from Pauli errors of a single type, such as $Z$ errors, are forced to come in pairs along one-dimensional strips of the lattice. This results from strips of stabilisers multiplying together to an operator that commutes with all $Z$ errors: these are one-dimensional symmetries of the stabiliser code under infinitely biased noise. We will refer to stabilisers in an ISG that are flipped by an error as anyons. In bias-tailored static codes, anyons can propagate within strips but cannot move outside the strip without changing Pauli type. We will see that in the case of the bias-tailored Floquet code, there exists a similar symmetry.
As in the CSS Floquet code, there are two disjoint decoding graphs, the A-type and B-type graphs, whose edges correspond to single-qubit Pauli errors. In even measurement subrounds, we perform B-type measurements and form detectors for the B-type graph while in odd measurement subrounds we only form detectors for the A-type graph. Note that under a more complicated noise model, such as one including measurement errors or two-qubit depolarising errors, edges can exist between the A- and B-type graphs and they are no longer disjoint.
@@ III.1 Persistent symmetry and two-dimensional decoding under biased noise
There is no constant stabiliser group when viewing the X3Z3 Floquet code as a subsystem code (in a subsystem code, the stabiliser group is defined as the center of the gauge group) [11, 10]. Surprisingly, despite this fact, there do exist operators at every time step that form a symmetry under infinitely $Z$-biased noise (without measurement errors). These symmetry operators at time step $t$, are formed by plaquette operators of a single type along vertical strips of the lattice; symmetries on unshaded and shaded strips are formed by A- and B-type plaquettes, respectively (see Fig. 2). We take, as these symmetries, the product of all such plaquette operators that commute with all operators in $\mathcal{M}_{t+1}$ (the set of check measurements to be performed in the next subround). To see why the symmetry is formed at each measurement subround, consider one particular time step as shown in Fig. 2(a), i.e., after red A-type measurements. As depicted, the product of red and blue A-type plaquettes is the identity on all qubits in the unshaded vertical strip, and acts as $Z$ on some qubits in the shaded vertical strips. A similar symmetry can be formed by blue and green B-type plaquettes on shaded strips in this subround.
@@
As a result of this symmetry, Pauli-$Z$ errors create anyons in pairs along vertical strips. An example of such anyons is shown in each measurement subround of Fig. 2(b). Note that the anyons change plaquette locations between subrounds, but always respect the symmetry. These anyons in Fig. 2 represent detectors that would be triggered if the Pauli-$Z$ error shown occurred at that time step. For example, if the $Z$ error occurs immediately after the red A-check measurements, it will trigger a red detector after subsequent blue A-check measurements and a blue detector after green A-check measurements.
@@
Unlike the symmetry in static stabiliser codes, this symmetry does not allow for one-dimensional decoding even without measurement errors. While in static codes, only measurement errors produce “time-like” edges, in Floquet codes, even Pauli errors produce time-like edges (between detectors formed at different times). To demonstrate this, we display in Fig. 3 a portion of the A-type detector graph under infinitely biased noise (with no measurement errors). As can be seen, there is a disconnected sub-graph defined along one vertical strip of the code lattice. Even considering only single-qubit Pauli noise without measurement errors, the decoding graph in the infinite bias regime is two-dimensional (i.e., the graph is planar). This results from the fact that neighbouring plaquettes are measured at different times. We emphasise that although the above discussion is based on $Z$ errors, the same analysis also holds for $X$ errors because $X$ and $Z$ operators are interchangeable for the X3Z3 code. As a result, the performance of the X3Z3 code under $X$-biased noise is expected to be the same as that under the $Z$-biased noise model studied in this paper.
@@ III.2 Bias-preserving parity-check circuits
While two-body measurements are native to certain architectures, for example, Majorana qubits [21, 48], most architectures require a syndrome extraction circuit to carry out such measurements. To maintain the high-performance of bias-tailored codes in these hardwares, the syndrome extraction circuits need to be constructed in a bias-preserving manner, such that they propagate errors in such a way so as to sustain the noise bias. To this end, we design two-qubit parity check circuits that preserve the $Z$ bias on data qubits. That is, the probability of $X$ and $Y$ errors on data qubits after these circuits is small (proportional to single-qubit $X$ or $Y$ error probabilities). A bias-preserving parity check measurement circuit can be constructed by using an ancilla circuit connecting two data qubits, which can be realised even in devices with minimal connectivity, such as the heavy-hexagonal layout. [45, 44].
@@
For a circuit to be generally bias-preserving, it needs to be constructed using only gates that do not change the error type under conjugation. For example, to preserve the $Z$-bias, $Z$ errors cannot be propagated to either $X$ or $Y$ errors on data qubits. One way to construct bias-preserving circuits is to use only CNOT gates in the measurement circuits as CNOT gates propagate errors to others with the same Pauli type. For the X3Z3 code, we need to construct bias-preserving circuits for three different kinds of parity checks: $XX$, $ZZ$ and $XZ$. All these three circuits can be constructed using only CNOT gates as shown in Fig. 4. As depicted, these circuits also include resets and measurements of the ancilla qubit in the Pauli bases, i.e., $Z$, $X$ and $Y$ bases for $ZZ$, $XX$ and $XZ$ checks, respectively. Note that the circuits are still bias-preserving even if the resets/measurements in the $X$ and $Y$ bases are compiled in terms of resets/measurements in the $Z$ basis with additional single-qubit Clifford gates. This is because the $Z$ error on any qubit at every time step in the compiled circuits still propagates to data qubits as a $Z$ error, or as a check operator about to be measured.
In realistic systems, errors can occur during the application of the gate. If conventional CNOT gates are used in the syndrome extraction circuits, then they may not preserve the noise bias. This is because a $Z$ error occuring on the target qubit during the application of the CNOT gate results in a combination of $Z$ and $X$ errors on the target qubit after the CNOT gate [49].
@@
One way to construct a circuit that preserves the noise bias in the presence of mid-gate errors is to use bias-preserving CNOT gates [49]. These bias preserving CNOT gates, however, are restricted to only specific platforms, such as the cat qubit platform. Here, we propose a general approach to constructing $Z$-bias preserving parity check circuits. While our proposed circuits can preserve only a specific type of noise bias, i.e., the $Z$ bias, these circuits are built using conventional two-qubit gates and thus can be implemented in many different architectures. The idea is to use the CZ gates as these do not propagate $Z$ errors, including the mid-gate $Z$ errors, into other kinds of errors [49]. If this is not possible, we can apply CNOT gates with ancillas as target qubits. In the following, we use this idea to construct fully bias-preserving circuits for the three parity checks shown in Fig. 4.
The simplest of these circuits is the depth-2 $ZZ$ measurement circuit, shown in Fig. 5(a) which is built using CZ gates. This circuit is equivalent to the one shown in Fig. 4. On the other hand, to implement an $XX$ measurement, we utilise a depth-4 circuit as shown in Fig. 5(b) which is constructed from CNOT gates which use ancillas as target qubits. Note that, in this circuit, the two single-qubit $X$ measurements on the data qubits in the middle of the circuit will not reveal the information of each of data qubit’s state since before the measurements, we entangle each of the data qubits to the ancilla qubit using a CNOT gate. The mid-circuit $X$ measurements on $\mathrm{Data}_{1}$ and $\mathrm{Data}_{2}$ qubits are equivalent to $X_{1}X_{a}$ and $X_{2}X_{a}$ measurements, respectively, where $X_{a}$ acts on the ancilla qubit in the |$\ket{0}$ state. The final $Z_{a}$ measurement disentangles the ancilla qubit, leaving the data qubits in an eigenstate of $X_{1}X_{2}$. This circuit therefore performs the required $XX$ check measurement.
@@
Finally, the $Z_{1}X_{2}$ checks can be measured by performing CZ gates directly between the data qubits, as shown in Fig. 5(c). However, without connections between the data qubits in hardware with a heavy-hexagonal lattice, these CZ gates have to be implemented using ancilla qubits in-between the data qubits. To this end, we prepare the ancilla in the $\ket{+}$ state and use a depth-3 circuit to implement this CZ gate (see the right-hand side of Fig. 5(c); this is adapted from a similar circuit in Ref. [44]). The end result is a depth-6 measurement circuit. In Fig. 5(c), we provide optional measurements, coloured in grey, which provide “flag information” for detecting $Z$ errors on the measurement ancilla that may have propagated to the data qubits.
@@
We emphasise that, even if the $X$ measurements in the parity-check circuits above are implemented using $Z$ measurements sandwiched by Hadamard gates, the circuits are still bias preserving since a $Z$ error occurring between those Hadamard gates is harmless. This is because a $Z$ error happening before the measurement is immediately absorbed by the $Z$ measurement without flipping its outcome, and a $Z$ error after the measurement is a stabiliser of the state (up to a sign), so does nothing. Moreover, if the Hadamard gates are noisy, the single-qubit gate errors in many architectures are not the predominant error and are usually much smaller than those of the two-qubit gate errors [50, 51, 52, 53].
@@
Figure 5: Measurement circuits for (a) $ZZ$, (b) $XX$, and (c) $Z_{1}X_{2}$ checks that preserve the $Z$ noise bias on data qubits. Data qubits are labelled by Data1 and Data2, while the middle qubit in each circuit is a measurement ancilla. In (c), the equality can be checked by noting that the depth-3 circuit before and after the measurement performs a next-nearest-neighbour CZ gate, if the measurement ancilla is initially in the $\ket{+}$ state. The grey measurements can be optionally included to provide flag information.
@@ IV Methods
We perform quantum memory experiment simulations of all codes mentioned above: the CSS and X3Z3 Floquet codes together with the P6 and XYZ2 honeycomb codes, with varying degrees of noise bias $\eta$. The noise bias $\eta=p_{z}/(p_{x}+p_{y})$ is defined as the ratio of the $Z$-error $p_{z}$ to other errors where the total physical error rate is $p=p_{x}+p_{y}+p_{z}$. As the noise asymmetry increases, the noise bias $\eta$ increases from $\eta=0.5$ which corresponds to a fully depolarising noise ($p_{x}=p_{y}=p_{z}=p/3$) to $\eta=\infty$ corresponding to a pure $Z$-bias noise ($p_{z}=p$, and $p_{x}=p_{y}=0$).
@@
We use two different biased noise models which are the generalisations of the standard code-capacity and standard depolarising entangling measurement (SDEM3) [9, 18] noise models to biased noise channels. The code-capacity noise model is one in which only single-qubit errors are applied to all data qubits at every measurement subround. We note that, since for Floquet codes one QEC round consists of several measurement subrounds, the single-qubit noise channel in the code-capacity model is applied several times in one QEC round instead of just once as is the case with static codes. The code-capacity noise model is often used as a preferred initial noise model to study before going to a more involved model as its simplicity often offers insight into understanding the code performance. On the other hand, the SDEM3 model is a more elaborate error model involving single-qubit noise channels after every single-qubit gate, measurement and reset, a two-qubit noise channel after every multi-Pauli product (MPP) parity measurement gate, and a classical flip after each measurement. As in Ref. [9], we assume that each of these error channels occurs with a total probability $p$. The SDEM3 noise model is therefore close to standard circuit-level noise and would be a more accurate description of a realistic noise channel, particularly in hardware with native two-qubit measurements, for example, Majorana qubits [48, 21]. To take into account biased noise for the MPP gates, we here generalise the SDEM3 noise model in Refs. [9, 18]. The details of the biased code-capacity and SDEM3 noise models can be found in Appendix B.
@@
We simulate the circuits and generate the error syndromes using Stim [54]. To decode the error syndromes, we apply a minimum-weight perfect matching (MWPM) decoder, which is implemented using PyMatching [47]. Our simulations are run for codes with effective distances $d_{\mathrm{eff}}\equiv d/2=4,6,8,10,12$, where the effective distance $d_{\mathrm{eff}}$ is defined as the minimum number of faults under SDEM3 depolarising noise that produce a logical error, which is half of the distance $d$ of the code-capacity noise model. The calculated effective distance differs between noise models since in the SDEM3 noise model, two-qubit errors occurring after an MPP parity gate count as a single fault, while under the code-capacity noise model they would count as two faults. We choose lattices of size $2d_{\mathrm{eff}}\times 3d_{\mathrm{eff}}$ for the distances mentioned, and with periodic boundary conditions. We simulate the memory experiments for $3d_{\mathrm{eff}}$ QEC rounds, with each round containing 6 subrounds of check measurements.

### v2

@@ III X3Z3 Bias-tailored Floquet Code
We now describe the X3Z3 bias-tailored Floquet code, which is shown in Figs. 1(b) and 1(c). This code is related to the CSS Floquet code by Hadamard gates applied to the qubits in alternating strips [i.e., the grey strips in Fig. 1(b)] along vertical non-trivial cycles of the lattice, and is a Floquetified version of the domain wall colour code [45]. As in the CSS Floquet code, the X3Z3 code also has two types of plaquettes and edges: one originates from the Pauli $X$ and the other from the Pauli $Z$ plaquettes and edges in the CSS code before the Hadamard deformation. We refer to these modified operators as A-type and B-type, respectively; these are shown in Fig. 1(b). The measurement sequence is analogous to the CSS Floquet code sequence: $r\text{A}\rightarrow g\text{B}\rightarrow b\text{A}\rightarrow r\text{B}\rightarrow g\text{A}\rightarrow b\text{B}$, where $c\text{A}$ represents the measurement of A-type check operators along $c$-coloured edges, and similarly for $c\text{B}$. Just as with the CSS Floquet code [13], this code can also be defined on a planar lattice with boundary.
The decoding of the X3Z3 Floquet code under $Z$-biased noise is simplified by the presence of a symmetry in the decoding graphs, which we will explain below. This is a space-time analogue of the symmetries present in bias-tailored static codes such as the XZZX code [41] and the domain wall colour code [45]. In such codes, under the infinitely-biased code-capacity noise model (with only single-qubit $Z$ errors, for example), syndromes are forced to come in pairs along one-dimensional (1D) strips of the lattice. This results from strips of stabilisers multiplying together to the identity along one domain and commuting with all $Z$ errors on all qubits: these are 1D symmetries of the stabiliser code under infinitely biased noise. We will refer to (plaquette) stabilisers in the ISG that are flipped by an error as anyons. In bias-tailored static codes, anyons can propagate within strips but cannot move outside the strip without changing their Pauli type. We will show that there also exists a similar symmetry in the bias-tailored Floquet code.
As in the CSS Floquet code, there are two disjoint decoding graphs for the X3Z3 Floquet code: the A-type and B-type graphs, whose edges correspond to single-qubit Pauli errors. In even measurement subrounds, we perform B-type measurements and form detectors for the B-type graph while in odd measurement subrounds we only form detectors for the A-type graph. Note that under a more complicated noise model, such as one that includes measurement or two-qubit errors, edges can exist between the A- and B-type graphs and they are no longer disjoint.
@@ III.1 Persistent symmetry and two-dimensional decoding under biased noise
There is no constant stabiliser group when viewing the X3Z3 Floquet code as a subsystem code (in a subsystem code, the stabiliser group is defined as the center of the gauge group) [13, 18]. Surprisingly, despite this fact, at every time step there do exist operators that form a symmetry under pure dephasing noise (without measurement errors). An example of such a symmetry is shown in Fig. 2 which is a symmetry on the unshaded strips formed by the A-type plaquettes. A similar symmetry can also be found on the shaded strips which are formed by B-type plaquettes. These symmetries, which are formed by plaquette operators of a single type along vertical strips of the lattice at time step $t$, also commute with all operators in $\mathcal{M}_{t+1}$ (the set of check measurements to be performed in the next subround). To see why the symmetries are formed at each measurement subround, consider one particular time step as shown in Fig. 2(b), i.e., after the red A-type measurements. As depicted, the product of red and blue A-type plaquettes is the identity on all qubits in the unshaded vertical strip, and yields Pauli $Z$ operators on some qubits in the neighbouring shaded vertical strips. At this subround, there is a similar symmetry that has the identity on the shaded strips, which is formed by the product of blue and green B-type plaquettes.
@@
As a result of each of these symmetries, Pauli-$Z$ errors create anyons in pairs along vertical strips. An example of such anyons is shown in each measurement subround of Fig. 2(a). Note that the anyons change plaquette locations between subrounds, but always respect the symmetry. The anyons depicted in Fig. 2 represent detectors that would be triggered if the Pauli-$Z$ error shown occurred at that time step. For example, if the $Z$ error occurs immediately after the red A-check measurements, it will trigger a red A detector after subsequent blue A-check measurements and a blue A detector after green A-check measurements.
@@
Unlike the symmetry in static stabiliser codes, this symmetry does not allow for 1D decoding even when measurement errors are absent. While in static codes, only measurement errors produce “time-like” edges, in Floquet codes, even Pauli errors produce time-like edges (between detectors formed at different times). To demonstrate this, we display in Fig. 3 a portion of the A-type detector graph under infinitely-biased single-qubit noise with no measurement errors. As can be seen, there is a disconnected subgraph defined along one strip of the code lattice. Even considering this simple noise model, the decoding graph in the infinite bias regime is two-dimensional (i.e., the graph is planar). This results from the fact that neighbouring plaquettes are measured at different times. We emphasise that although the above discussion is based on $Z$ errors, the same analysis also holds for $X$ errors because $X$ and $Z$ operators are interchangeable for the X3Z3 code. As a result, the performance of the X3Z3 code under $X$-biased noise is expected to be the same as that under the $Z$-biased noise model.
@@ III.2 Bias-preserving parity-check circuits
While two-body measurements are native to certain architectures, for example, Majorana [10, 56, 57] and photonic [58, 17] platforms, most architectures require quantum circuits to carry out such measurements. To maintain the high-performance of bias-tailored codes in these hardwares, the parity-check measurement circuits need to be constructed in a bias-preserving manner, such that they do not propagate frequently-occurring errors to rarely-occurring errors. To this end, we design two-qubit parity check circuits that preserve the $Z$ bias on data qubits. That is, the probability of $X$ and $Y$ errors on data qubits after these circuits remains small (proportional to single-qubit $X$ or $Y$ error probabilities). Our bias-preserving parity check measurement circuits areconstructed by using an ancilla circuit connecting two data qubits, which can be realised even in devices with minimal connectivity, such as the heavy-hexagonal layout [53, 52].
@@
For a circuit to be bias preserving, it can be constructed using only gates that do not change the error type under conjugation. For example, one way to construct bias-preserving circuits is to use only CNOT gates in the measurement circuits as they only propagate errors to others with the same Pauli type. For the X3Z3 code, we need to construct bias-preserving circuits for three different kinds of parity checks: $ZZ$, $XX$ and $ZX$. All these three circuits can be constructed using only CNOT gates as shown in Fig. 4. As depicted, these circuits also include resets and measurements of the ancilla qubit in the Pauli bases, i.e., $Z$, $X$ and $Y$ bases for $ZZ$, $XX$ and $ZX$ checks, respectively. Note that the circuits are still bias-preserving even if the resets/measurements in the $X$ and $Y$ bases are compiled in terms of resets/measurements in the $Z$ basis with additional single-qubit Clifford gates. This is because a $Z$ error on any qubit at every time step in the compiled circuits still propagates to data qubits as a $Z$ error, or as a check operator about to be measured.
@@
The parity-check circuits shown in Fig. 4, however, may not preserve the noise bias if errors happen during the CNOT gates. This is because a $Z$ error occurring on the target qubit during the application of a conventional (non-bias preserving) CNOT gate may result in a combination of $Z$ and $X$ errors on the target qubit after the CNOT gate [59]. One way to remedy this problem is to use bias-preserving CNOT gates [59]. These gates, however, are available only in specific platforms such as cat qubits. Here, we propose a more general approach to constructing $Z$-bias preserving parity check circuits without requiring the bias-preserving gates. While our proposed circuits can preserve only a specific type of noise bias, i.e., the $Z$ bias, these circuits are built using conventional (which may not be bias-preserving) two-qubit gates and thus can be implemented in many different architectures.
Our proposed fully bias-preserving parity-check circuits are depicted in Fig. 5. These circuits are constructed using either CZ gates or CNOT gates targeting ancilla qubits, which only propagate Pauli errors on the ancilla qubits as $Z$ errors on the data qubits. While mid-gate $Z$ errors may result in $X$ errors on the ancilla qubits, such errors do not degrade the bias of errors affecting the data qubits. We note that while the $ZZ$ parity check measurements are performed by reading out the ancilla qubits like in standard syndrome extraction circuits, the $XX$ and $ZX$ parity values are obtained by measuring the data qubits after some two-qubit gates are applied on them. These two qubits gates are required for the parity measurements so as not to reveal each individual data qubit’s state. For example, the two mid-circuit $X$-measurements in Fig. 5(b), are equivalent to reading out the $X_{1}X_{a}$ and $X_{2}X_{a}$ parity values, where $X_{a}$ acts on the ancilla qubit which is prepared in the |$\ket{0}$ state. The final $Z_{a}$ measurement disentangles the ancilla qubit, leaving the data qubits in an eigenstate of $X_{1}X_{2}$. The circuit therefore performs the required $XX$ check measurement.
@@
The mixed-type $Z_{1}X_{2}$ parity circuit can be constructed using CZ gates between the data qubits, as shown in Fig. 5(c). If the connections between the data qubits are not available in the device such as in hardwares with a heavy-hexagonal lattice, these CZ gates have to be implemented using ancilla qubits in-between the data qubits. We show such an implementation of next-nearest-neighbour CZ gates in the right-hand side of Fig. 5(c) (this is adapted from Ref. [52]). In the figure, we also present optional measurements, coloured in grey, which provide “flag information” for detecting $Z$ errors on the measurement ancilla that may have propagated to the data qubits.
@@
We emphasise that, even if the $X$ measurements in the parity-check circuits above are implemented using $Z$ measurements sandwiched by Hadamard gates, the circuits are still bias preserving since a $Z$ error occurring between those Hadamard gates is harmless. This is because a $Z$ error happening before the measurement is immediately absorbed by the $Z$ measurement without flipping its outcome, and a $Z$ error after the measurement is a stabiliser of the state (up to a sign), so does nothing. Moreover, the circuit will not change the noise bias significantly even if the Hadamard gates are noisy. The reason is because in many architectures, the single-qubit gate errors are not the predominant error and are usually much smaller than the two-qubit gate errors [60, 61, 62, 63].
@@
Figure 5: Measurement circuits for (a) $ZZ$, (b) $XX$, and (c) $Z_{1}X_{2}$ checks that preserve the $Z$ noise bias on data qubits. Data qubits are labelled by Data1 and Data2, while the middle qubit in each circuit is a measurement ancilla. In (c), the equality can be checked by noting that the depth-3 circuit before and after the mid-circuit measurement performs a next-nearest-neighbour CZ gate, if the measurement ancilla is initially in the $\ket{+}$ state. The grey measurements can be optionally included to provide flag information.
@@ IV Methods
To study the performance of our proposed X3Z3 code, we perform quantum memory experiment simulations, with varying degrees of noise bias $\eta$, of this code along with three other codes: CSS, P6 and XYZ2 Floquet codes. The noise bias $\eta=p_{Z}/(p_{X}+p_{Y})$ is defined as the ratio of the $Z$-error $p_{Z}$ to other errors where the total physical error rate is $p=p_{X}+p_{Y}+p_{Z}$. As the noise asymmetry increases, the noise bias $\eta$ increases from $\eta=0.5$, which corresponds to fully depolarising noise ($p_{X}=p_{Y}=p_{Z}=p/3$), to $\eta=\infty$, corresponding to a pure $Z$-biased noise ($p_{Z}=p$, and $p_{X}=p_{Y}=0$).
@@
We use two different biased noise models which are the generalisations of the standard code-capacity and standard depolarising entangling measurement (SDEM3) [9, 15] noise models to biased noise channels. The code-capacity noise model is one in which only single-qubit errors are applied to all data qubits at every measurement subround. We note that, since for Floquet codes one QEC round consists of several measurement subrounds, the single-qubit noise channel in the code-capacity model is applied several times in one QEC round instead of just once as is the case with static codes. The code-capacity noise model is often used as a preferred initial noise model to study before going to a more involved model as its simplicity often offers insight into understanding the code performance.
On the other hand, the SDEM3 model is a more elaborate error model involving single-qubit noise channels after every single-qubit gate, measurement and reset, a two-qubit noise channel after every multi-Pauli product (MPP) parity measurement gate, and a classical flip after each measurement. As in Ref. [9], we assume that each of these error channels occurs with a total probability $p$. The SDEM3 noise model is therefore close to standard circuit-level noise and would be a more accurate description of a realistic noise channel, particularly in hardwares with native two-qubit measurements, for example, Majorana [56, 10, 57] and photonic [58, 17] architectures. To take into account biased noise for the MPP gates, we here generalise the SDEM3 noise model in Refs. [9, 15]. The details of the biased code-capacity and SDEM3 noise models can be found in Appendix B.
@@
We simulate the circuits and generate the error syndromes using Stim [64] (example Stim circuits and threshold data are provided at Ref. [65]). To decode the error syndromes, we apply a minimum-weight perfect matching (MWPM) decoder, which is implemented using PyMatching [55]. Our simulations are run for codes with effective distances $d_{\mathrm{eff}}\equiv d/2=4,6,8,10,12$, where the effective distance $d_{\mathrm{eff}}$ is defined as the minimum number of faults under SDEM3 depolarising noise that produce a logical error. This is half of the distance $d$ of the code-capacity noise model. The calculated effective distance differs between noise models since in the SDEM3 noise model, two-qubit errors occurring after an MPP parity gate count as a single fault, while under the code-capacity noise model they would count as two faults. We choose lattices with periodic boundary conditions and sizes $L\times 3L/2$. Here, $L=d=2d_{\mathrm{eff}}$ is the lattice length which varies with respect to the code distance. We simulate the memory experiments for $3d_{\mathrm{eff}}$ QEC rounds, with each round containing 6 subrounds of check measurements.